It is known that 66% of all planted seeds germinates. What is the probability that after sowing 2386 seeds, the number of germinated seeds will be in the range of (1569, 1587)? That more than 1587 seeds will germinate?

Answer by Boreal(15235)

mean=np=1574.76
variance=np(1-p)=535.5184, multiply top by 0.34
sd=sqrt(V)=23.14
lower end is using continuity correction factor z=(1568.5-1574.76)/23.14=-0.27
and z=(1587.5-1574.76)23.14=0.55
probability z is in those ranges is 0.3153
-
more than 1587 is z>0.55 and that probability is 0.2912